OpenStudy (anonymous):

Can someone explain to me how to rewrite this equation in vertex form?? y = x^2 – 6x + 4

OpenStudy (anonymous):

It's basically completing the square. Can you do that or do you still need help?

OpenStudy (anonymous):

I still need help. It's in my brain somewhere but can you just take me through it step by step? :/

OpenStudy (anonymous):

I'd be glad to. y = x^2 – 6x + 4 is in the form y = ax^2 + bx + c where a=1, b=-6, and c=4. We catch a big break here because a=1 and that greatly simplifies things. I'll show you how to work these when a=1. If you get another problem with a=some other number, I can show you that later if you like. I'm going to work with a, b, and c and you can substitute so you'll be doing something as well.

OpenStudy (anonymous):

y=(a(x-h)^2) + k is vertex form. The vertex is (h,k). +a is smiley, -a is frowney - that tells you if the parabola opens up or down. To convert an equation to vertex form complete the square. y=x^2+6x+4 Step 1: move the loose number (the constant term) to the y side. y-4=x^2 + 6x Step 2: Factor out whatever multiplies x^2. Here it's 1, so this step does nothing except place the right side in ()'s. y-4=(x^2 + 6x) Step 3: Make a space after the constant on the y side, place the factor before x^2 before that space and the sign before that space is the same as the sign of the number in front of x^2. Again, since it's 1, no need for that part. y-4+()=(x^2 + 6x) Take 1/2 the coefficient of the x term (the number in front of x). Remember its sign. Square it. Add it to both sides inside the (). y-4+(9)=(x^2 + 6x +9)

OpenStudy (anonymous):

So, we have y = x^2 + bx + c -> y = [x^2 + bx + (b/2)^2] + c - (b/2)^2

OpenStudy (anonymous):

y = [x^2 + (b/2)]^2 + c - (b/2)^2 and now you have your vertex form. This is the general solution for when a=1 and can be used as a pattern for any problem of that type now. Vertex at (-[b/2], c - [b/2]^2). It's a parabola opening up and that vertex is a minimum value.

OpenStudy (anonymous):

Is this starting to make sense to you?

OpenStudy (anonymous):

Uhhh, I was following @Grim0519 better... No offence, I'm a visual lerner. and all the words confuse me sometimes..

OpenStudy (anonymous):

You can now take that equation from my fourth post and substitute a=1, b=-6, and c=4 and then you're done!

OpenStudy (anonymous):

np. Whatever way you learn better is what matters.

OpenStudy (anonymous):

teamwork i like that...

OpenStudy (anonymous):

I still don't get it. Where do I put them? so far I have y-4(9)=(x^2+6x+9) .... Where do I go from there?

OpenStudy (anonymous):

Multiply out the "a times the squared coefficient" part on the left-hand side (remember, in this one a = 1 so that does nothing), and convert the right-hand side to squared form. (This is where you use that sign you kept track of earlier, putting that sign in the middle of the squared expression.) y-4+(9)=(x+3)^2 Simplify - combine like terms. y+5=(x+3)^2 Move the constant term from the right back to the left. y=(x+3)^2 - 5 Write in vertex form y=(a(x-h)^2) + k. In other words if the squared term is x+h write it as x-(-h). If the k term is negative, write it as + (-k). y=(x-(-3))^2 + (-5) Now the values for h and k are clear.

OpenStudy (anonymous):

Okk, so the answer is y=(x-(-3))^2+(-5) ?

OpenStudy (anonymous):

y=x^2+6x+4 =x^2+6x +4 =x^2+6x+9+4-9 complete the square by dividing 6 by 2 and squaring the result. =(x+3)^2-5 you should have ended with this.

OpenStudy (anonymous):

Ok got it. Thank you! Would you be willing to help me with one more?

OpenStudy (anonymous):

sure shoot...

OpenStudy (anonymous):

Same thing, converting to vertex form. y=x^2+14x+50 ......

OpenStudy (anonymous):

gimme a sec...

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

@Grim0519 Shouldn't that last problem solution be y = (x - 3)^2 - 5 instead of y = (x + 3)^2 - 5 ? Maybe you had a typo somewhere?

OpenStudy (anonymous):

f(x) = a(x-h)^2 + k "a" represents the leading coefficient: a = 1 (h,k) are the coordinate of the vertex of parabola. We'll create a perfect square within the given quadratic: f(x) = (x^2 - 14x + 49) + 1 f(x) = (x - 7)^2 + 1 The coordinates of the vertex of the parabola are (7 , 1). The vertex form of the given quadratic f(x) = x^2 - 14x + 50 is f(x) = (x - 7)^2 + 1.

OpenStudy (anonymous):

i didnt think so...let me recheck my work...

OpenStudy (anonymous):

So the answer to the second one is y=(x-7)^2+1 ?

OpenStudy (anonymous):

Or f(x) rather. not y.

OpenStudy (anonymous):

yes ma'am

OpenStudy (anonymous):

Yay! Thank you soooo much! Both of you!

OpenStudy (anonymous):

Wait, that one is wrong also.

OpenStudy (anonymous):

The second one is y=(x+7)^2+1 and the first one is y = (x - 3)^2 - 5. Grim, you're flipping the sign around. Check again.

OpenStudy (anonymous):

Oh dear....

OpenStudy (anonymous):

I fixed it. Thanks for catching that tcarroll! You're a life saver.

OpenStudy (anonymous):

sorry about that the second one was wrong......

OpenStudy (anonymous):

The confirmation is in the x^1 term. For that first equation, you have to end up with -6x. You won't get that with (x + 3)^2. The first one was wrong also.

OpenStudy (anonymous):

the first also thanks @tcarroll010

OpenStudy (anonymous):

np. Good working with both of you. Good effort all around.

OpenStudy (anonymous):

Olivia, no disrespect to Grim, because he's a good hard worker. Really a good hard worker, but his method is a bit off. You might want to review my method to get the right answer. It doesn't matter to me whose method you follow, but I want you to get the right answer for these problems in the future.

OpenStudy (anonymous):

definitely...